# Empowering Multi-Robot Cooperation via Sequential World Models

### v1

@@ 5.1 Comparisons
Figure 4: Performance comparisons of SeqWM with other baselines. Task in Bi-DexHands report the episode return, while Multi-Quad (gray background) reports success rate. Bold lines indicate the mean over multiple seeds, with shaded regions denoting the 95% confidence intervals.
Results on Bi-DexHands. We select six representative tasks, including object transfer tasks (Over,CatchAbreast,CatchOver2Underarm), which require the two hands to transfer an object under different relative positions and grasping postures; and functional manipulation tasks (BottleCap,Pen,Scissors), which involve precise bimanual operations to achieve specific functional goals, such as opening a bottle cap, removing a pen lid, or spreading a pair of scissors.
@@
Results on Multi-Quad. The Multi-Quad tasks require two quadruped robots to cooperate under spatial constraints and dynamic interactions. In Gate, the robots are required to pass through a narrow gate as quickly as possible without collision 666 We extend Gate to 3–5 agents in Appendix C.1 to evaluate SeqWM’s scalability.. In PushBox, they jointly push a large box to a designated target location. In Shepherd,, the two quadruped robots (as sheepdogs) cooperatively guide another robot (as sheep) to a target area (as sheep pen).
In Gate and Shepherd, it rapidly approaches near-100% success rates within the early phase, significantly surpassing baselines in terms of sample efficiency. In PushBox, SeqWM is the only method capable of completing the task, whereas all baselines fail to push the box to the target. This superior performance stems from SeqWM’s sequential world model, which enables each agent to plan actions conditioned on its predecessors’ intentions, thereby enhancing coordination.
@@ 5.2 Empowered Cooperative Behaviors
We further visualize the behaviors learned by SeqWM, showing that it not only learns stable policies in high-dimensional state and action spaces, but also achieves advanced cooperative behaviors, including predictive adaptation, role division, temporal alignment, and intention sharing.
Bi-DexHands Behaviors. Trajectory visualizations in Figure 5 show that SeqWM acquires sophisticated coordination strategies. In Catch-Over2Underarm, the receiving hand anticipates the object’s motion based on predictive communication from its partner and pre-adapts its grasp posture, enabling a precise and stable catch near the target. In Pen, both hands achieve near-perfect temporal alignment, grasping the pen body and cap almost simultaneously to complete the removal action efficiently. These behaviors stem from SeqWM’s combination of world model–based planning and sequential decision-making. The leading agent first predicts and plans its actions, explicitly transmitting future trajectories to its partner. This intention sharing enables the receiving agent to act on foresight rather than delayed perception, producing synchronized, task-specific coordination and enhancing collaborative efficiency.
@@
Figure 5: Trajectory visualizations of the tasks selected for Bi-DexHands with SeqWM.
@@
Multi-Quad Behaviors. Figure 6 illustrates the coordinated behavior learned by SeqWM in PushBox. From $t=1$ (frame A) to $t=2$ (frame B), the quadrupeds navigate to opposite sides of the box, establishing an effective pushing configuration. Between $t=2$ (frame B) and $t=4$ (frame D), both maintain high positive $x$-axis velocities, indicating sustained forward pushing force. At $t=3$ (frame C), Robot 2 produces a downward $y$-axis velocity, adjusting the push direction toward the target, while Robot 1 gradually increases its negative $y$-axis velocity to assist with directional control. As the box nears the target, Robot 1 reduces its $x$-axis velocity to avoid overshooting. These behaviors demonstrate that SeqWM supports not only effective force coordination for object propulsion but also fine-grained directional adjustments, resulting in precise and efficient task completion.
@@
Figure 6: Behavior visualizations in PushBox. The first row shows the execution process, where the box is significantly larger than the robots, requiring coordinated efforts from both quadrupeds to complete the task. The second row visualizes the trajectories of the robots and the box, with the right side showing the x-axis and y-axis velocities and orientations of each robot.
@@ 5.3 Real-World Deployment
Real-World Results. The real-world experimental setup is detailed in Appendix C.5, and the results are shown in Figure 7.
Figure 7: Real-world results of multi-robot cooperation experiments. The trajectories of Robot 1, Robot 2, and the Sheep are marked in different colors.
In PushBox, the two quadrupeds approach the box from opposite sides and coordinate their pushing forces and directions to move it toward the target. Between frames (D-F), Robot 1 moves forward to provide the main pushing force, while Robot 2 makes slight lateral adjustments to steer the box. The overall pushing pattern, including the division of roles and the gradual directional adjustments, closely matches the behavior observed in simulation, confirming a successful sim-to-real transfer.
In Gate, two clear yielding events are observed. Between frames (C-D), Robot 1 slows down and waits for Robot 2 to pass first, demonstrating priority management in constrained spaces. After crossing (frames E-F), Robot 2 actively veers aside to leave sufficient space for Robot 1, enabling smooth passage without collisions. These behaviors reflect SeqWM’s trajectory prediction and intention-sharing capabilities, allowing natural, efficient yielding without high-frequency communication.
In Shepherd, Robot 1 accelerates between frames (A-B), causing the Sheep to move left. To prevent the Sheep from hitting the left gate frame, Robot 1 retreats while Robot 2 advances between frames (C-D). This maneuver drives the Sheep away from Robot 2 and into the target area. The sequence highlights SeqWM’s capacity for predictive coordination and adaptive role allocation, where the one agent’s motion influences the sheep robot’s response and the another agent adapts accordingly to achieve the common goal.
@@ 5.4 Ablation Studies
Sequential Sample Generation. To evaluate the contribution of the sequential world model in SeqWM, we replace it with centralized and decentralized architectures, ensuring all models have an equal number of parameters for a fair comparison. Using BottleCap, we collect 50K environment steps with random actions and train each model for 2.5K steps using the loss in Eq. (3). After training, we gather 1K additional steps to measure dynamics and reward prediction errors across different horizons. As shown in Figure 8, the sequential and centralized models achieve similarly low errors, both substantially outperforming the decentralized model. The results confirm the advantage of sequential prediction, where each agent conditions its output on the predictions of its predecessors, yielding more accurate and coherent rollouts.
@@
Communication Function. We replaced the concat communication function in SeqWM with alternative fusion mechanisms, including MLP, cross-attn, and RNN, and evaluated them on the BottleCap task. The results in Figure 9 show that the simplest concat approach achieves the highest and most stable performance. This advantage stems from two factors: (i) concat preserves the complete communication content, allowing the dynamics and reward predictors to autonomously identify and exploit the most informative features during training; and (ii) it introduces no additional learnable parameters, thereby maintaining stable gradient propagation in long-horizon prediction. Moreover, we observe that RNN-based fusion even underperforms the no-communication baseline (dec), which we attribute to its sensitivity to input ordering—an undesirable property in multi-agent communication scenarios lacking a fixed semantic sequence.
@@ 5.5 Additional Experiments
    Visualizations the learned behaviors on all other tasks in Appendix C.2;
@@
    An ablation analysis of the actor to illustrate its complementary role with the planner in Appendix C.3;
@@
    An analysis of the planner’s time cost and early-stopping heuristic in Appendix C.4;
@@
    An evaluation of SeqWM’s scalability with respect to the number of agents in Appendix C.1.
@@ 6 Conclusion
This paper presented SeqWM, a novel model-based MARL framework that integrates the sequential paradigm into world model learning and planning. By structurally decomposing joint dynamics into autoregressive, agent-wise models, SeqWM offers a principled approach that reduces modeling complexity, lowers communication overhead, and naturally enables intention sharing through predicted trajectories. This methodological innovation not only improves scalability but also facilitates the emergence of advanced cooperative behaviors such as predictive adaptation, temporal alignment, and role division. Extensive experiments in Bi-DexHands and Multi-Quad show that SeqWM achieves state-of-the-art performance with superior sample efficiency, while real-world deployment on quadruped robots confirms that these cooperative behaviors transfer reliably from simulation to physical platforms. Beyond empirical results, SeqWM demonstrates that sequential paradigms provide a scalable and theoretically grounded principle for structuring multi-agent cooperation, paving the way for more robust and efficient deployment of cooperative policies in real-world systems.
@@ A.2 Low-Pass Action Smoothing
Figure 11: Visualization of the low-pass filter. (Left): Amplitude–frequency response of the low-pass filter. (Middle): Filtering effects on random signals at 20% cutoff ratios, the different colors represent different action dimensions. (Right): Effects of low-pass filtering on control commands in the PushBox task, with different colors representing different agents.
@@ B.1 Pseudocode
Figure 12: Overview of implementation details. (Left): SEM Normalization. (Middle): Symlog transformation, where the $y$-axis represents the values encoded by two-hot encoding, and the $x$-axis represents the values aligned with the scalar target. (Right): Normalized $Q$-scale and return in the Dex-BottleCap task.
@@ B.2 Model Architecture
The proposed sequential world model is composed of five components: encoder, dynamics predictor, reward predictor, critic, and actor, all implemented using MLPs. We report the network configurations and the number of parameters in each module on the Dex-BottleCap task $\left(|\mathcal{O}|=221,|\mathcal{A}|=26\right)$ in Table 2.
@@ B.3 Reward and Value Prediction
Discrete Regression. The reward prediction and critic in SeqWM are modeled as discrete regression problems, which helps enhance robustness to extreme values. Take reward prediction for example, we discretize the reward range into $B$ intervals and use two-hot encoding (Bellemare et al., 2017; Hafner et al., 2025) to map the scalar reward to a $B$-dimensional vector. The two-hot encoding assigns weights to the two adjacent bins, reducing quantization errors and providing smoother gradients. Formally, for a scalar reward signal $r$, its two-hot encoding is defined as:
@@
Percentile Scaling of Critic
We employ a percentile-based scaling strategy to stabilize Q-value magnitudes. Specifically, we computes the dynamic range between the 5-th and 95-th percentiles of each batch, which effectively suppresses the influence of outliers. This range is updated smoothly using an EMA (Exponential Moving Average) controlled by a factor $\tau$:
@@ Appendix C Additional Experiments
C.1 Scalability with the Number of Agents
Multi-Agent MuJoCo. To evaluate the scalability of SeqWM with respect to the number of agents, we conduct supplementary experiments on the 6a-Cheetah task in the MA-Mujoco (Multi-Agent MuJoCo) (Peng et al., 2021) environment. MA-Mujoco partitions a robot into multiple agents according to different body parts; in the 6a-Cheetah task, six agents are required to coordinate to control a single joint to run faster. To achieve scalability, we modify the communication protocol in Eq. (2) to transmit only the action sequence:
@@ C.1 Scalability with the Number of Agents
Gate with More Robots. We also evaluate scenarios with more robots on the Gate task, with success rates reported in Figure 14. As the number of robots increases, the task becomes more complex, but SeqWM still successfully learns effective cooperative strategies, achieving nearly 100% success rates in all settings within 4M time steps. Additionally, we visualize the learned behaviors of SeqWM with different numbers of robots in Figure 15, Figure 16, and Figure 17.
@@
In the 5-robot Gate task shown in Figure 17, the robots exhibit similar coordinated behaviors. Across agents, the X Command traces show pronounced wave-like undulations: troughs denote the robot waiting at the gate, whereas crests indicate the robot currently traversing the passage. Around $t{\approx}3$ (Frame C), an implicit queue emerges in the “X Command” panel: Robot 2 sustains a near-unity forward command while Robot 1 and Robot 3 reduce theirs moderately; Robot 4 and Robot 5 exhibit the deepest dips, preventing congestion at the bottleneck. From $t{=}3{\to}4$ (Frame D), Robot2 clears the gate first and issues a negative $y$-command to drift downward and vacate the corridor (“Y Command”), with yaw traces staying close to zero for all robots (“Robots’ Yaw”). Next, Robot 1 increases its $x$-command to pass second while slightly offsetting upward in $y$ to avoid interference. Finally, from $t{=}4{\to}5$ (Frame E), Robot 3 proceeds third, followed by Robot 4 and Robot 5, which had maintained the lowest $x$-commands during the queuing phase—completing a smooth, collision-free sequence (order: 2 $\rightarrow$ 1 $\rightarrow$ 3 $\rightarrow$ 4 $\rightarrow$ 5).
C.2 Additional Behavior Visualization

### v2

@@ 5.1 Comparisons
Figure 3: Performance comparisons on selected tasks of SeqWM with other baselines. Task in Bi-DexHands report the episode return, while Multi-Quad (gray background) reports success rate. Bold lines indicate the mean over multiple seeds, with shaded regions denoting the 95% confidence intervals. The results on all other tasks are reported in Figure 13 in Appendix C.1.
Results on Bi-DexHands. The representative tasks in Bi-DexHands include object transfer tasks (Over,CatchAbreast,CatchOver2Underarm), which require the two hands to transfer an object under different relative positions and grasping postures; and functional manipulation tasks (BottleCap,Pen,Scissors), which involve precise bimanual operations to achieve specific functional goals, such as opening a bottle cap, removing a pen lid, or spreading a pair of scissors.
@@
Results on Multi-Quad. In Gate, the robots are required to pass through a narrow gate as quickly as possible without collision. In PushBox, they jointly push a large box to a designated target location. In Shepherd,, the two quadruped robots (as sheepdogs) cooperatively guide another robot (as sheep) to a target area (as sheep pen).
In Gate and Shepherd, it rapidly approaches near-100% success rates within the early phase, significantly surpassing baselines in terms of sample efficiency. In PushBox, SeqWM is the only method capable of completing the task, whereas all baselines fail to push the box to the target. 333 SeqWM-MARL shows the behavior comparison between SeqWM and the next-best baseline on BottleCap and PushBox. This superior performance stems from SeqWM’s sequential world model, which enables each agent to plan actions conditioned on its predecessors’ intentions, thereby enhancing coordination.
@@ 5.2 Empowered Cooperative Behaviors
We further visualize the behaviors learned by SeqWM, showing that it not only acquires stable policies in high-dimensional state and action spaces, but also achieves advanced cooperative behaviors, including predictive adaptation, temporal alignment, and role division.
Bi-DexHands Behaviors. In Catch-Over2Underarm, the throwing hand first performs prediction and planning, explicitly transmitting future trajectories to the catching hand. Guided by this message, the catching hand then exhibits predictive adaptation by anticipating the object’s motion and landing point and proactively adjusting its grasping posture. As shown in Frames C–D, the catching hand lowers and opens in advance, aligning its posture with the predicted landing point to enable a reliable grasp. In Pen, two hands achieve near-perfect temporal alignment by exchanging predictions of future actions in advance. As a result, they grasp the pen body and cap almost simultaneously in Frame D and efficiently complete the extraction in Frames E–F, substantially enhancing cooperative efficiency.
@@
Figure 4: Trajectory visualizations of Catch-Over2Underarm and Pen with SeqWM.
@@
Multi-Quad Behaviors. Figure 5 further shows the role division learned by SeqWM in the Multi-Quad-PushBox. In Frames A–B ($t=1\to 2$), the two quadruped robots navigate to opposite sides of the box, establishing an effective pushing configuration. In Frames B–D ($t=2\to 4$),, both maintain high positive $x$-axis velocities, indicating continuous forward pushing force. At Frame C ($t=3$), Robot 2 produces a downward $y$-axis velocity, adjusting the push direction toward the target, while Robot 1 gradually increases its negative $y$-axis velocity to assist in directional control. As the box approaches the target, Robot 1 reduces its $x$-axis velocity to avoid overshooting. These behaviors demonstrate that SeqWM supports not only effective force coordination but also fine-grained directional adjustments, resulting in precise and efficient task completion.
@@
Figure 5: Behavior visualizations in PushBox. The first row shows the execution process, where the box is significantly larger than the robots, requiring coordinated efforts from both quadrupeds to complete the task. The left side of second row visualizes the trajectories of the robots and the box, with the right side showing the x-axis and y-axis velocities and orientations of each robot.
5.3 Scalability to More Agents
We extend the Gate to 3–5 agents to evaluate the scalability of SeqWM with respect to the number of agents, and the behavioral visualizations of 5-robot-Gate are presented in Figure 6.
Figure 6: Visualization of the learned behaviors on 5-robot-Gate.
As the robots approach the narrow gate, they exhibit predictive adaptation, with certain agents proactively decelerating or waiting to avoid potential congestion. For instance, at Frame B ($t{\approx}2$), Robot 3 maintains a near-unity positive x-command, while the other robots moderately reduce their forward commands. In terms of temporal alignment, the x-command trajectories reveal a clear wave-like alternation, where the peak sequence (3 $\rightarrow$ 2 $\rightarrow$ 1 $\rightarrow$ 4 $\rightarrow$ 5) mirrors the actual passing order, reflecting a dynamic “first-pass–then-follow” sequence. Overall, the team establishes a coordinated rhythm of “prediction–waiting–passing–yielding,” which enables efficient multi-robot traversal under constrained environmental conditions. Further quantitative analyses and visualizations on 3--5 robot Gate and Mujoco-6a-Cheetah are provided in Appendix C.2.
@@ 5.4 Real-World Deployment
The real-world experimental setup is detailed in Appendix C.6, and the results are shown in Figure 7.
Figure 7: Real-world results of multi-robot cooperation tasks. The trajectories of Robot 1, Robot 2, and the Sheep are marked in different colors.
In PushBox, the two quadrupeds approach the box from opposite sides and coordinate their pushing forces and directions to move it toward the target. Between Frames D-F, Robot 1 moves forward to provide the main pushing force, while Robot 2 makes slight lateral adjustments to steer the box. The overall pushing pattern, including the division of roles and the gradual directional adjustments, closely matches the behavior observed in simulation, confirming a successful sim-to-real transfer.
In Gate, two clear yielding events are observed. Between Frames C-D, Robot 1 slows down and waits for Robot 2 to pass first, demonstrating priority management in constrained spaces. After crossing Frames E-F, Robot 2 actively veers aside to leave sufficient space for Robot 1, enabling smooth passage without collisions. These behaviors reflect SeqWM’s trajectory prediction and intention-sharing capabilities, allowing natural, efficient yielding without high-frequency communication.
In Shepherd, Robot 1 accelerates between Frames A-B, causing the Sheep to move left. To prevent the Sheep from hitting the left gate frame, Robot 1 retreats while Robot 2 advances between Frames C-D. This maneuver drives the Sheep away from Robot 2 and into the target area. The sequence highlights SeqWM’s capacity for predictive coordination and adaptive role allocation, where the one agent’s motion influences the sheep robot’s response and the another agent adapts accordingly to achieve the common goal.
@@ 5.5 Ablation Studies
Sequential Sample Generation. To evaluate the contribution of the sequential paradigm in SeqWM’s world model, we replace it with centralized and decentralized architectures, ensuring all models have an equal number of parameters for a fair comparison. Using BottleCap, we collect 50K environment steps with random actions and train each model for 2.5K steps using the loss in Eq. (3). After training, we gather 1K additional steps to measure dynamics and reward prediction errors across different horizons. As shown in Figure 8, the sequential and centralized models achieve similarly low errors, both substantially outperforming the decentralized model. The results confirm the advantage of sequential prediction, where each agent conditions its output on the predictions of its predecessors, yielding more accurate and coherent rollouts.
@@
Communication Function. We replaced the concat communication function in SeqWM with alternative fusion mechanisms, including MLP, cross-attn, and RNN, and evaluated them on BottleCap. The results in Figure 9 show that the simplest concat approach achieves the highest and most stable performance. This advantage stems from two factors: (i) concat preserves the complete communication content, allowing the dynamics and reward predictors to autonomously identify and exploit the most informative features during training; and (ii) it introduces no additional learnable parameters, thereby maintaining stable gradient propagation in long-horizon prediction. Moreover, we observe that RNN-based fusion even underperforms the no-communication baseline (dec), which we attribute to its sensitivity to input ordering—an undesirable property in multi-agent communication scenarios lacking a fixed semantic sequence.
@@ 5.6 Additional Experiments
    Visualizations of the learned behaviors on all other tasks in Appendix C.3 and SeqWM-MARL;
@@
    An ablation analysis of the complementary roles of the actor and planner in Appendix C.4;
@@
    An evaluation of SeqWM’s scalability with respect to the number of agents in Appendix C.2;
@@
    An experimental analysis of time consumption and early stopping in the planner in Appendix C.5.
@@ 6 Conclusion
This paper presented SeqWM, a novel model-based MARL framework that integrates the sequential paradigm into world model learning and planning. By structurally decomposing joint dynamics into autoregressive, agent-wise models, SeqWM offers a principled approach that reduces modeling complexity and naturally enables intention sharing through predicted trajectories. This methodological innovation not only improves scalability but also facilitates the emergence of advanced cooperative behaviors such as predictive adaptation, temporal alignment, and role division. Extensive experiments in Bi-DexHands and Multi-Quad show that SeqWM achieves state-of-the-art performance with superior sample efficiency, while real-world deployment on quadruped robots confirms that these cooperative behaviors transfer reliably from simulation to physical platforms. Beyond empirical results, SeqWM demonstrates that sequential paradigms provide an efficient and scalable principle for structuring multi-agent cooperation, paving the way for more robust and efficient deployment of cooperation in physical multi-robot systems.
Future Work. Benefiting from the integration of the sequential paradigm and agent-wise world models, SeqWM naturally extends to heterogeneous robot teams and human–robot semantic understanding. With each agent maintaining an independent world model, the framework accommodates diverse dynamics and sensing modalities, enabling cooperation among quadrupeds, manipulators, and aerial robots. Moreover, the explicit trajectory rollouts can be shared not only across robots but also with humans as interpretable intention signals, fostering transparent collaboration, mutual understanding, and trust in human–robot teams.
Ethics Statement
This work focuses on advancing multi-robot cooperation through model-based reinforcement learning. All experiments were conducted in simulated environments and on standard quadruped robot platforms, with no involvement of humans, animals, or sensitive personal data. The proposed methodology focuses on technical contributions to the fields of reinforcement learning, multi-agent systems, and robotics, without ethical implications beyond standard academic research practices.
Reproducibility Statement
To ensure the reproducibility of the results, we provide open-source code at https://github.com/zhaozijie2022/seqwm-marl. Hyperparameters and training procedures are detailed in Appendix B. All baseline comparisons use publicly available implementations, with the documented parameter settings as referenced in the respective sections.
Use of Large Language Models
Large Language Models (LLMs) were used in the preparation of this paper exclusively for writing assistance and language polishing. The conceptualization of the research, methodology design, experimental implementation, and analysis were all conducted entirely by the authors. The authors take full responsibility for the content of this paper.
@@ A.2 Low-Pass Action Smoothing
Figure 11: Visualization of the low-pass filter. (Left): Amplitude–frequency response of the low-pass filter. (Middle): Filtering effects on random signals at 20% cutoff ratios, the different colors represent different action dimensions. (Right): Effects of low-pass filtering on control commands in PushBox, with different colors representing different agents.
@@ B.1 Pseudocode
Figure 12: Overview of implementation details. (Left): SEM Normalization. (Middle): Symlog transformation, where the $y$-axis represents the values encoded by two-hot encoding, and the $x$-axis represents the values aligned with the scalar target. (Right): Normalized $Q$-scale and return in Dex-BottleCap.
@@ B.2 Model Architecture
The proposed sequential world model is composed of five components: encoder, dynamics predictor, reward predictor, critic, and actor, all implemented using MLPs. We report the network configurations and the number of parameters in each module on Dex-BottleCap $\left(|\mathcal{O}|=221,|\mathcal{A}|=26\right)$ in Table 1.
@@ B.3 Reward and Value Prediction
Discrete Regression. The reward prediction and critic in SeqWM are modeled as discrete regression problems, which helps enhance robustness to extreme values. Take reward prediction for example, we discretize the reward range into $B$ intervals and use two-hot encoding (Bellemare et al., 2017; Hafner et al., 2025) to map the scalar reward to a $B$-dimensional vector. The two-hot encoding assigns weights to the two adjacent bins, reducing quantization errors and providing smoother gradients (Zhu et al., 2025b). Formally, for a scalar reward signal $r$, its two-hot encoding is defined as:
@@
Percentile Scaling of Critic We employ a percentile-based scaling strategy to stabilize Q-value magnitudes. Specifically, we computes the dynamic range between the 5-th and 95-th percentiles of each batch, which effectively suppresses the influence of outliers. This range is updated smoothly using an EMA (Exponential Moving Average) controlled by a factor $\tau$:
@@ Appendix C Additional Experiments
C.1 Comparisons on Other Tasks
We report additional comparison results on other tasks to complement Figure 3.
Figure 13: Comparison results on other tasks.
C.2 Scalability with the Number of Agents
Multi-Agent MuJoCo. To evaluate the scalability of SeqWM with respect to the number of agents, we conduct supplementary experiments on the 6a-Cheetah in the MA-Mujoco (Multi-Agent MuJoCo) (Peng et al., 2021) environment. MA-Mujoco partitions a robot into multiple agents according to different body parts; in 6a-Cheetah, six agents are required to coordinate to control a single joint to run faster. To achieve scalability, we modify the communication protocol in Eq. (2) to transmit only the action sequence:
@@ C.2 Scalability with the Number of Agents
Gate with More Robots. We also evaluate scenarios with more robots on Gate, with success rates reported in Figure 15. As the number of robots increases, the task becomes more complex, but SeqWM still successfully learns effective cooperative strategies, achieving nearly 100% success rates in all settings within 4M time steps. Additionally, we visualize the learned behaviors of SeqWM with different numbers of robots in Figure 16, Figure 17, and Figure 18.
@@
Across agents, the X Command traces show pronounced wave-like undulations: troughs denote the robot waiting at the gate, whereas crests indicate the robot currently traversing the passage. Around Frame B ($t{\approx}2$), an implicit queue emerges in the “X Command” panel: Robot 3 sustains a near-unity forward command while Robot 1 and Robot 2 reduce theirs moderately; Robot 4 and Robot 5 exhibit the deepest dips, preventing congestion at the bottleneck. In Frames (C-D) ($t{=}3{\to}4$), Robot 2 clears the gate first and issues a negative $y$-command to drift, vacating space for others. Robot 1 then accelerates forward to pass second, slightly shifting upward in $y$ to prevent interference, followed by Robot 3. Finally, Robot 4 and Robot 5, which had consistently yielded earlier, complete the crossing. This progression underscores role specialization, with some agents acting as initiators and others as supporters, while the group achieves a well-aligned, collision-free sequence (order: 3 $\rightarrow$ 2 $\rightarrow$ 1 $\rightarrow$ 4 $\rightarrow$ 5).
C.3 Additional Behavior Visualization